Prove the following please simplify and then and then answer....cot x - tan x = 2 cot 2x

Grade:11

1 Answers

Arun
6 years ago
cot x - tan x = cosx/sinx - sinx/cosx = (cos^2 x - sin^2 x)/(sinxcosx) = cos2x/[(1/2)sin2x] = 2cot 2x